OpenStudy (anonymous):

A pair of equations is shown below: y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (anonymous):

@Johnvdb

OpenStudy (anonymous):

Haha already? Alright so, very similar to the last question:

OpenStudy (anonymous):

Thats just to show you the answer. Now if we are going to use substitution and elimination

OpenStudy (anonymous):

oksy

OpenStudy (anonymous):

Alright i will draw you some things for elimination, and tell me if you don't get it

OpenStudy (anonymous):

Actually nah i don't like elimination, substitution is always easier. so, because both of these equations are equal to y, we set them equal to each other: see how 8x - 9 = y, well we can just put in 8x - 9 = 4x - 1 then we solve: 8x - 9 = 4x - 1 4x = 8 x = 2

OpenStudy (anonymous):

The reason we can plug in the equations and set them equal to each other is because they both are equal to y, and this is very helpful for the future. And the answer to part 2 is in the attachment I sent earlier

OpenStudy (anonymous):

ohh!

OpenStudy (anonymous):

Now if you need me to repeat anything at all, let me know :)

OpenStudy (anonymous):

well where exactly do they intersect

OpenStudy (anonymous):

Alright now it's your turn :) pull up desmos.com/calculator

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

you're going to go on the left side and type the equation into one box, then click right underneath that box and type the next equation in and it will graph it for you, and you will see where they intersect

OpenStudy (anonymous):

1.7?

OpenStudy (anonymous):

screenshot what you're doing and send it as an attachment so I can help you with what you've got going wrong for the future

OpenStudy (anonymous):

can screen shot on this P.C

OpenStudy (anonymous):

darn ok, well let me give you a quick tutorial then

OpenStudy (anonymous):

That's the right one, ok so all you have to do is click right in that box and type, and the graph will pop up. You can either type in 8x - 9 or y = 8x - 9 (the calculator knows what you're saying either way)

OpenStudy (anonymous):

now right underneath that box, you can click and another box will be outlined and let you type in it

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

They intersect at (2,7)

OpenStudy (anonymous):

YES THATS WHAT I GOT!!!!!!!!

OpenStudy (anonymous):

There you go :) Now I completely believe in your ability to use this tool in the future, it can solve so many of your questions, but if you still aren't sure just tag me and I will do my best to help you :)

OpenStudy (anonymous):

THANKS APRECIATE IT

OpenStudy (anonymous):

Now go forth and do great things!